OpenStudy (anonymous):

Hi again, I have to simplfy the following expression from (A-(A/3))/4sqrt(A/3). The answer is 1/2 x sqrt(A/3), but I'm unsure how to arrive at this simplification.

OpenStudy (anonymous):

a-a/3=2a/3 2a/3/4sqrt(a/3) here, 2/4=1/2 a/3/sqrt(a/3)=sqrt(a/3) so u get it :)

OpenStudy (anonymous):

got it thanks!!!!